### v1

# Efficient Nearest Neighbor based Uncertainty Estimation for Natural Language Processing Tasks

Abstract
Trustworthy prediction in Deep Neural Networks (DNNs), including Pre-trained Language Models (PLMs) is important for safety-critical applications in the real world. However, DNNs often suffer from uncertainty estimation, such as miscalibration. In particular, approaches that require multiple stochastic inference can mitigate this problem, but the expensive cost of inference makes them impractical. In this study, we propose -Nearest Neighbor Uncertainty Estimation (NN-UE), which is an uncertainty estimation method that uses the distances from the neighbors and label-existence ratio of neighbors. Experiments on sentiment analysis, natural language inference, and named entity recognition show that our proposed method outperforms the baselines or recent density-based methods in confidence calibration, selective prediction, and out-of-distribution detection. Moreover, our analyses indicate that introducing dimension reduction or approximate nearest neighbor search inspired by recent NN-LM studies reduces the inference overhead without significantly degrading estimation performance when combined them appropriately.
Efficient Nearest Neighbor based Uncertainty Estimation for Natural Language Processing Tasks
Wataru Hashimoto, Hidetaka Kamigaito, Taro Watanabe Nara Institute of Science and Technology {hashimoto.wataru.hq3, kamigaito.h, taro}@is.naist.jp
1 Introduction
In order to use Deep Neural Networks (DNNs) including Pre-trained Language Models in safety-critical regions, uncertainty estimation is important. By improving the predictive uncertainty, the prediction will be calibrated Guo et al. (2017),111”Calibration” means the confidence of the model aligns with its accuracy. or improve selective prediction performance, which is predictive performance when there is a choice to abstain from model prediction Galil et al. (2023). On the other hand, DNNs often fail to quantify the predictive uncertainty, for example, causing miscalibrated prediction Guo et al. (2017). Such uncertainty estimation performance problems can be mitigated by the PLMs, such as BERT Devlin et al. (2019) or DeBERTa He et al. (2021b), that are self-trained on large amounts of data Ulmer et al. (2022), although, there is still a need for improvement Desai and Durrett (2020).
To solve the problem of uncertainty estimation, multiple probabilistic inferences such as MC Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017) are generally effective. On the other hand, these methods require multiple stochastic inferences for a single data instance, which leads to high computational cost, and makes them impractical for real world application. To obtain reasonable predictive uncertainty without multiple inferences, Temperature Scaling Guo et al. (2017) is generally used, which scales logits with a temperature parameter. Furthermore, density-based methods such as Density Softmax Bui and Liu (2024a) and Density Aware Calibration (DAC) Tomani et al. (2023), which correct the model outputs based on estimated density, have achieved promising very recent years in terms of uncertainty estimation performance and inference cost. However, both Density Softmax and DAC only uses the density of training data. Therefore, we can see that these methods only capture epistemic uncertainty that comes from the knowledge of the model. To improve the uncertainty estimation performance, we also need to consider aleatoric uncertainty that comes from the variance of the data Hüllermeier and Waegeman (2019).
In this study, we propose -Nearest Neighbor Uncertainty Estimation (NN-UE), a new density-based uncertainty estimation method that does not require multiple inferences. NN-UE uses the labels of neighbors obtained from NN search to correct the confidence as illustrated in Figure 1. Our method weights logits according to the distance between the input example and its neighbors in the datastore created by the training data. In addition, the logit weights are further biased by the ratio of the model’s predicted label matched with the labels in neighbors. As a result, our method requires only a single forward inference of the model.
First, our experiments show that NN-UE improves the uncertainty estimation performance of existing baselines in sentiment analysis, natural language inference, and named entity recognition in both in-domain and out-of-domain settings by combining neighbor label information and distances from neighbors. Second, we show that the smaller the number of instances of neighbor search in NN-UE, the better the uncertainty performance tends to be slight. Finally, to solve the latency in NN-UE for token-level tasks, such as sequence-labeling based name entity recognition, we show that approximate nearest neighbor search or dimension reduction in NN-UE improves the inference speed without degrading the uncertainty estimation performance much more, while combining them leads to degrading the uncertainty performance. Our code will be available after acceptance.
2 Related Work
Uncertainty Estimation
Many prior studies on uncertainty estimation focus primarily on image classification. One approach for improved uncertainty estimation is the sampling-based approach of performing multiple stochastic predictions for one data instance. The approach includes MC Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017). However, their practical application is limited due to the computationally expensive inference. As an alternative to improve uncertainty, the sampling-free approaches do not demand multiple stochastic inferences, e.g., Temperature Scaling Guo et al. (2017), Deep Deterministic Uncertainty Mukhoti et al. (2023), and Posterior Network Charpentier et al. (2020). Very recently, the density-based methods Bui and Liu (2024a); Tomani et al. (2023); Bui and Liu (2024b) show good performance on latency and uncertainty estimation.
-Nearest Neighbor Language Models / Machine Translation
-Nearest Neighbor Language Model (NN-LM) Khandelwal et al. (2020) has been proposed, which performs linear interpolation of NN probability based on distance from neighbors and base model probability, in the language modeling task. -Nearest Neighbor Machine Translation (NN-MT) applied the NN-LM framework to machine translation Khandelwal et al. (2021). NN-LM and NN-MT have been successful because they enhance predictive performance through the memorization and use of rich token representations of pre-trained language models and mitigate problems such as a sparsity comes from low-frequency tokens Zhu et al. (2023). The main issue on NN-LM and NN-MT is the inference overhead, and there are several studies to solve this problem. He et al. (2021a) employs datastore compression, adaptive retrieval, and dimension reduction to reduce computational overhead with retaining perplexity. Deguchi et al. (2023) dramatically improves decoding speed by dynamically narrowing down the search area based on the source sentence.
Uncertainty Estimation for Natural Language Processing Tasks
Studies about uncertainty estimation for NLP tasks are limited when compared with those for image datasets. Kotelevskii et al. (2022) has shown excellent performance in classification with rejection tasks and out-of-distribution detection tasks using uncertainty scores using density estimation results. Vazhentsev et al. (2022) performed misclassification detection using Determinantal point processes, spectral normalization, and loss regularization in text classification and NER. Vazhentsev et al. (2023) proposed the new uncertainty method which combined epistemic and aleatoric uncertainty estimation methods for text classification tasks. In this study, we comprehensively investigate the density-based methods such as Density Softmax and DAC in NLP tasks, inspired by the success of memorization and use in NLP region Zheng and Jiang (2022), such as NN-LM and NN-MT. In addition, our study attempts to improve the uncertainty estimation performance by making more effective use of memorized examples, such as introducing label information of neighbor examples.
3 Base Methods
In this section, we describe the density-based methods for uncertainty estimation. We use the following notations: denotes the logits for class , denotes the calibrated probability for class , means the dimension of the final output layer, and denotes the number of classes.
3.1 Density Softmax
Density Softmax Bui and Liu (2024a) obtains confidence by weighting logits with normalized log-likelihood from a trained density estimator. In this study, we use RealNVP Dinh et al. (2017) as the density estimator (details for the density estimator are in Appendix A). is the parameters of the density estimator; is the representation obtained from the input example ; is the normalized log-likelihood from the density estimator, then the corrected confidence is written as
In Density Softmax, the closer the normalized log-likelihood to zero, the closer the prediction to Uniform distribution Bui and Liu (2024a). Density Softmax achieves reasonable latency and competitive uncertainty performance with state-of-the-art methods at the cost of demanding the density estimator training and multiple base model training.
3.2 Density Aware Calibration (DAC)
DAC Tomani et al. (2023) scales the logits by using sample-dependent temperature
| where | ||||
are the weights for every layer of the base model, is the averaged distance from NN search on -th layer, and is the bias term. are optimized using the L-BFGS-B method Liu and Nocedal (1989) based on the loss in the validation set. In the original DAC paper, the uncertainty estimation performance tends to improve with the increase in the number of layer’s representation Tomani et al. (2023). Therefore, we report the performance when all the hidden representations in each layer of the base PLMs are used.
DAC avoids the problem of the high training cost in Density Softmax. However, both DAC and Density Softmax use only the density of the training data representations for uncertainty estimation. Considering that the uncertainty is mainly composed of epistemic uncertainty and aleatoric uncertainty, DAC and Density Softmax only represent the epistemic uncertainty because they use only the distance from neighbors or the data likelihood. Therefore, the uncertainty estimates could be further improved by introducing the concept of aleatoric uncertainty.
4 Proposed Method: -Nearest Neighbor Uncertainty Estimation (NN-UE)
Our proposed method, NN-UE, is based on the intuition that a model should give high confidence only when the predicted instance exists in the region with high density of training data and a high proportion of the model’s predicted label exists in the neighbors sharing the same label, as illustrated in Figure 1. To represent this intuition, we introduce a scoring method based on NN and labels of the neighbors in the following formula:
| where | ||||
is the number of neighbors from NN search, is the number of examples where the predicted label and the label of the neighbors are same, is the distance between the -th representation obtained by NN search and the representations of training data.222Note that NN-UE is also ”accuracy-preserving” same as DAC because is a scalar, not a class-wise score. The parameters , , and are optimized using the L-BFGS-B method based on the loss in the validation set. As with Density Softmax, the closer is to zero, the closer the prediction is to Uniform distribution, which allows us to better estimate confidence of the prediction. Unlike DAC, which relies only on the distances to neighbors and Density Softmax, which uses the log-likelihood from the density estimator, the label term in NN-UE increases the uncertainty when there is the high aleatoric uncertainty region that corresponds to the high variation of labels in neighbors of the predicted query representation.
5 Experimental Settings
| Tasks | Datasets | Train | Dev | Test | |
| SA | IMDb | 2 | 25,000 | 12,500 | 12,500 |
| Yelp | 2 | - | - | 19,000 | |
| NLI | MNLI | 3 | 392,702 | 4,907 | 4,908 |
| SNLI | 3 | - | - | 9,824 | |
| NER | OntoNotes 5.0 () | 37 | 10,683 | 1,295 | 1,357 |
| OntoNotes 5.0 () | 37 | - | - | 2,327 | |
| OntoNotes 5.0 () | 37 | - | - | 1,366 |
5.1 Tasks and Datasets
We measure the uncertainty estimation performance on Sentiment Analysis (SA), Natural Language Inference (NLI), and Named Entity Recognition (NER) in In-domain (ID) and Out-of-Domain settings. Dataset statistics are described in Table 1.
Sentiment Analysis (SA)
Natural Language Inference (NLI)
Named Entity Recognition (NER)
extracts the named entities, such as a person, organization, or location. The NER task was carried out in the framework of sequence labeling. We regard the OntoNotes 5.0 dataset Pradhan et al. (2013) broadcast news () domain as ID, and newswire () and telephone conversation () domains as OOD.
5.2 Existing Methods
Softmax Response (SR)
is a trivial baseline, which treats the maximum score from output of the base model’s softmax layer as the confidence Cordella et al. (1995).
Temperature Scaling (TS)
is a calibration technique by which the logits (i.e., the outputs of the final layer before the softmax) are divided by a temperature parameter before applying the softmax function Guo et al. (2017). We optimized by L-BFGS on dev set loss.
Label Smoothing (LS)
Monte-Carlo Dropout (MC Dropout)
is a regularization technique that be used for estimating uncertainty in neural networks Gal and Ghahramani (2016). This approach involves executing the model times with varying dropout masks and computing the mean softmax output across these runs. In our experiments, we set for all evaluations, and the dropout rate is 0.1.
5.3 Training Setting
In all experiments, we train and evaluate the models on a single NVIDIA A100 GPU with 40GB of memory. We used 333microsoft/deberta-v3-base and 444microsoft/mdeberta-v3-base He et al. (2023), as the transformer encoder from Wolf et al. (2020) pre-trained model checkpoints. Cross-entropy loss is minimized by AdamW Loshchilov and Hutter (2019) with a linear scheduler Goyal et al. (2017). The batch size is 32, and gradient clipping is applied with the maximum norm of 1. The initial learning rate was set to 1e-5. All experiments are run five times, and we report the mean and standard deviation of the scores.
| Methods | IMDb (In-domain) | Yelp (Out-of-domain) | ||||||
| ECE (↓) | MCE (↓) | AUROC (↑) | E-AURC (↓) | ECE (↓) | MCE (↓) | AUROC (↑) | E-AURC (↓) | |
| SR | 4.420.41 | 24.063.52 | 98.350.10 | 10.602.81 | 4.691.20 | 21.026.74 | 98.150.39 | 11.843.15 |
| TS | 4.100.31 | 20.435.01 | 98.450.21 | 11.362.82 | 5.101.19 | 19.701.35 | 98.200.46 | 12.914.12 |
| LS | 1.880.41 | 21.504.53 | 98.360.45 | 14.527.24 | 2.530.43 | 16.473.51 | 98.300.45 | 12.906.09 |
| MC Dropout | 4.280.27 | 23.743.52 | 98.570.12 | 9.171.74 | 4.330.54 | 20.172.79 | 98.280.25 | 10.012.01 |
| Density Softmax | 4.230.36 | 27.106.92 | 98.340.08 | 11.392.48 | 4.990.48 | 21.983.68 | 98.090.24 | 13.052.72 |
| DAC | 1.510.33 | 14.172.73 | 98.360.37 | 12.726.15 | 2.350.12 | 6.442.23 | 97.860.60 | 14.265.90 |
| NN-UE (w/o label) | 1.330.36 | 13.133.24 | 98.650.13 | 9.360.36 | 2.230.29 | 6.332.76 | 98.270.11 | 10.970.91 |
| NN-UE | 0.950.12 | 9.021.39 | 98.640.12 | 7.970.61 | 1.450.15 | 4.171.52 | 98.230.39 | 9.920.61 |
| Methods | MNLI (In-domain) | SNLI (Out-of-domain) | ||||||
| ECE (↓) | MCE (↓) | AUROC (↑) | E-AURC (↓) | ECE (↓) | MCE (↓) | AUROC (↑) | E-AURC (↓) | |
| SR | 8.360.61 | 37.617.53 | 97.030.12 | 31.292.23 | 9.770.55 | 36.6114.05 | 96.070.17 | 37.620.67 |
| TS | 2.731.86 | 15.8111.05 | 97.060.02 | 31.241.86 | 3.921.79 | 18.1310.69 | 96.080.13 | 38.402.06 |
| LS | 2.890.14 | 28.647.90 | 96.560.55 | 37.9812.64 | 3.970.45 | 23.186.17 | 95.610.40 | 44.189.18 |
| MC Dropout | 8.130.65 | 30.176.83 | 96.970.06 | 32.312.25 | 9.620.53 | 28.905.03 | 96.100.11 | 37.192.99 |
| Density Softmax | 8.420.43 | 36.205.78 | 97.030.10 | 32.563.29 | 10.090.40 | 33.594.57 | 95.960.19 | 41.432.25 |
| DAC | 1.420.30 | 18.7910.81 | 96.920.10 | 33.892.60 | 2.270.16 | 11.553.48 | 96.080.07 | 40.233.00 |
| NN-UE (w/o label) | 1.280.43 | 16.5311.45 | 97.090.10 | 30.222.80 | 2.120.36 | 10.006.07 | 96.120.16 | 37.334.70 |
| NN-UE | 1.410.47 | 10.772.34 | 97.180.09 | 23.831.29 | 1.800.37 | 5.121.47 | 96.000.22 | 34.972.48 |
Datastore Construction
It is necessary to preserve the representation of the data for training a density estimator in Density Softmax and NN search in DAC and NN-UE. We maintain final layer representations corresponding to CLS tokens in SA and NLI. In NER, we stored the hidden representation of the final layer as a token representation corresponding to the beginning of the word.
-Nearest Neighbor Search
5.4 Evaluation Metrics
To evaluate the confidence calibration performance, we choose Expected Calibration Error (ECE) and Maximum Calibration Error (MCE). For selective prediction, we evaluate Area Under the Receiver Operator Characteristic curve (AUROC) and Excess-Area Under the Risk-Coverage curve (E-AURC). Evaluation metrics computation details are described in Appendix B.
6 Results
6.1 Sentiment Analysis
We present the uncertainty estimation results for sentiment analysis. Table 2 shows the results of in-domain and out-of-domain uncertainty estimation. NN-UE consistently outperforms existing methods in terms of ECE, MCE, and E-AURC. In AUROC, LS outperforms in OOD setting, but NN-UE outperforms existing methods in ID setting. Furthermore, the proposed method clearly outperforms DAC that uses an ensemble of neighbor search results for each hidden representation, by adding the label term.
6.2 Natural Language Inference
In NLI, we evaluate the uncertainty estimation performance (calibration and selective prediction) and the out-of-distribution detection performance.
6.2.1 Uncertainty Estimation
First, we show the results of in-distribution and out-of-distribution uncertainty estimation in the NLI task using the DeBERTaV3 model in Table 3. Similar to Section 6.1, NN-UE shows a significant improvement in ECE, MCE and E-AURC. Furthermore, in AUROC, NN-UE outperforms existing methods in both ID and OOD settings. Galil et al. (2023) have reported that improving calibration performance does not necessarily lead to improving selective prediction performance, but our proposed method improves both. On the other hand, the degree of improvement is greater for calibration performance. Specifically, the largest improvement is obtained on SNLI, where NN-UE reduces MCE by more than 31.49 % pt compared to SR.
| Methods | FPR@95 (↓) | AUROC (↑) | AUPR-In (↑) | AUPR-Out (↑) |
| SR | 93.752.68 | 52.200.91 | 68.110.26 | 35.762.03 |
| TS | 91.012.01 | 52.701.08 | 68.080.79 | 37.771.81 |
| LS | 94.213.09 | 52.361.08 | 68.500.54 | 35.382.17 |
| MC Dropout | 93.261.71 | 52.681.26 | 68.210.81 | 36.411.63 |
| Density Softmax | 93.362.24 | 51.220.78 | 67.220.24 | 35.461.93 |
| DAC | 92.951.63 | 52.781.45 | 68.511.01 | 36.381.94 |
| NN-UE (w/o label) | 91.331.21 | 53.220.74 | 68.510.63 | 37.961.18 |
| NN-UE | 88.491.37 | 54.142.01 | 69.021.07 | 40.222.19 |
6.2.2 Out-of-Distribution Detection
Following the previous study Tomani et al. (2023), we carried out the experiments in the out-of-distribution detection task. Out-of-distribution detection is the task that determines whether the data is in-domain or not. This task is based on the intuition that we want to return predictions with high confidence in ID but with low confidence in predictions in OOD. We evaluated the out-of-distribution detection performance by using maximum softmax probability as the uncertainty score, and report FPR@95 (the FPR when the TPR is 95%), AUROC, Area Under the Precision-Recall curve (AUPR)-in and AUPR-out. AUPR-in indicates the AUPR score when ID samples are treated as positive; AUPR-out is vice versa.
| Methods | (In-domain) | (Out-of-domain) | (Out-of-domain) | ||||||
| ECE (↓) | MCE (↓) | E-AURC (↓) | ECE (↓) | MCE (↓) | E-AURC (↓) | ECE (↓) | MCE (↓) | E-AURC (↓) | |
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 17.050.69 | 37.063.13 | 81.494.17 | 21.202.03 | 42.605.84 | 76.055.72 |
| TS | 5.340.43 | 75.7121.96 | 19.631.22 | 12.760.62 | 26.573.97 | 72.904.72 | 19.690.95 | 47.727.34 | 71.878.83 |
| LS | 6.460.74 | 50.9926.73 | 24.931.19 | 14.780.61 | 30.542.84 | 81.506.98 | 20.992.16 | 65.4017.16 | 76.657.33 |
| MC Dropout | 6.760.64 | 53.1326.07 | 19.913.39 | 15.271.01 | 33.604.93 | 77.213.72 | 21.931.63 | 56.5612.32 | 75.689.30 |
| Density Softmax | 7.320.25 | 59.0527.76 | 25.172.63 | 16.100.62 | 44.6621.67 | 80.148.50 | 24.401.84 | 62.5010.46 | 80.066.27 |
| DAC | 1.620.42 | 42.9628.25 | 21.472.90 | 7.910.75 | 25.285.15 | 75.242.43 | 14.421.57 | 47.9220.98 | 80.728.19 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 8.780.62 | 24.911.81 | 70.104.03 | 14.610.67 | 35.267.16 | 65.418.11 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 7.500.42 | 16.532.61 | 74.275.43 | 14.150.33 | 39.846.02 | 71.819.04 |
Table 4 shows the out-of-distribution detection results when using MNLI/SNLI datasets as ID/OOD, respectively, in model. NN-UE consistently shows the out-of-distribution detection performance improvement.
6.3 Named Entity Recognition
In order to evaluate NLP tasks other than simple multi-class classification, we evaluate the effectiveness of the proposed method for uncertainty estimation in NER. Since NER focuses on entities, it is necessary to obtain the confidence of the entity. In this research, we use the product of the confidence of the tokens that construct the entity as the confidence level of the entity.
Table 5 shows the results of in-distribution and out-of-distribution uncertainty estimation using the OntoNote 5.0 dataset in the mDeBERTaV3 model. The proposed method improves the uncertainty estimation performance in all cases except ECE of DAC in the and domains, often resulting in large improvements compared to the SR. Moreover, we can see that TS and MC Dropout, which are often effective in multi-class classification, are not effective in evaluating entities like NER. On the other hand, E-AURC in NER is consistently better without using the kNN-UE label term.
| Text | As long as you go into this movie with the understanding that it’s not going to contain any historical fact whatsoever, it’s not bad.<br /><br />It’s on par with Sam Raimi’s Ḧercules: The Legendary Journeys,̈ as far as plot, acting, humour, and production values are concerned. You’ll see the similarities at several points. Most of the fight scenes are not as good however and the film suffers from that. … |
| Label | negative |
| SR & NN-UE pred. | positive |
| SR conf. | 0.76 |
| NN-UE (w/o label) conf. | 0.71 |
| NN-UE conf. | 0.60 |
| 11 |
6.4 Case Study: Effects of the Label Term in NN-UE for a Misclassified Example
Table 6 shows SR and NN-UE confidences, and in NN-UE for a misclassified example. In this case, SR and NN-UE make incorrect prediction even though the true label is negative. However, the confidence is appropriately reduced by including the distances from the neighbors in NN-UE, compared to SR. Moreover, by using the information that there are only 11 examples in neighbors with the same label as the predicted label among the neighbors obtained by NN search, our NN-UE shows that the confidence is further reduced.
7 Analyzes
7.1 Impact of Top-
To understand the behavior of NN-UE, we evaluated the performance in uncertainty estimation when changing the number of neighbors during NN execution.
Figure 2 shows the results for SA, and Figure 3 shows the results for NER. As is noticeable in NER, the smaller , the better uncertainty estimation tends to be. Since our method averages the distance to the top examples, logits are scaled to be more limited to neighbors by reducing . It is assumed that the uncertainty estimation performance is slightly improved as the NN-UE scoring becomes more dependent on neighbor data if is small.
| Methods | SNLI | OntoNotes 5.0 |
| SR | 21.590.76 | 5.750.27 |
| TS | 21.640.07 | 5.790.17 |
| LS | 21.700.07 | 5.800.19 |
| MC Dropout | 396.861.10 | 101.980.83 |
| Density Softmax | 22.020.05 | 6.020.07 |
| DAC | 2346.6236.06 | 326.001.41 |
| NN-UE (w/o label) | 23.020.04 | 10.360.21 |
| NN-UE | 23.070.05 | 10.480.12 |
| OntoNotes 5.0 (In-domain) | OntoNotes 5.0 (Out-of-domain) | |||||||
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 2.490.08 | 17.050.69 | 37.063.13 | 81.494.17 | 5.750.27 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 4.940.10 | 8.780.62 | 24.911.81 | 70.104.03 | 10.360.21 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| +PQ () | 1.960.31 | 31.3318.74 | 20.231.27 | 3.320.05 | 7.570.45 | 16.432.73 | 74.385.36 | 7.230.16 |
| +Clustering () | 1.920.31 | 28.5511.24 | 20.131.22 | 3.310.06 | 7.600.41 | 17.122.35 | 74.345.35 | 7.330.21 |
| +DR () | 2.140.37 | 33.5210.84 | 20.121.26 | 2.870.04 | 8.080.53 | 24.035.46 | 74.505.42 | 6.200.20 |
| Only DR () | 1.800.36 | 27.8513.80 | 20.131.29 | 3.410.10 | 7.540.45 | 16.422.73 | 74.305.44 | 7.750.24 |
| Methods | OntoNotes 5.0 | OntoNotes 5.0 |
| NN-UE | 100.0 | 100.0 |
| +PQ () | 21.30 | 51.68 |
| +Clustering () | 18.60 | 11.04 |
| +DR () | 0.02 | 0.04 |
| Only DR () | 43.98 | 20.35 |
7.2 Impact of Efficient Nearest Neighbor Search Techniques
In this subsection, we investigate the inference time and uncertainty estimation performance when applying approximate nearest neighbor search techniques and dimension reduction when executing NN search in NN-UE. As shown in Table 7, in the sequence labeling based NER that requires the NN search execution per token, it takes twice as much inference time as SR. On the other hand, in NN-LM Khandelwal et al. (2020), dimension reduction and approximate NN search techniques are effective to improve inference speed while maintaining perplexity He et al. (2021a); Xu et al. (2023). Therefore, inspired by these works for faster NN-LM, we investigate how the approximate nearest neighbor search techniques, such as Product Quantization Jégou et al. (2011) or clustering, and dimension reduction affect the uncertainty estimation and inference speed of our proposed method: NN-UE.
Product Quantization
Product Quantization (PQ) Jégou et al. (2011) is a data compression technique based on vector quantization. In PQ, a -dimensional representation is divided into subvectors and quantized by performing -means clustering on the vectors in each subspace. Vector quantization can significantly reduce the amount of memory occupied by vectors.555For example, raw datastore in kNN-UE is 636MB on OntoNotes 5.0 , but PQ reduces it to 10MB. In addition, by calculating the distance between compressed PQ codes, we can efficiently calculate the estimated value of the original Euclidean distance.
Clustering
The original NN-LM uses an inverted file index (IVF) technique that speeds up the search by dividing the representation into clusters by -means and searching for neighbors based on centroids. In this study, we evaluate the uncertainty estimation performance and inference speed when the number of clusters .
Dimension Reduction
In general, Transformer-based models such as PLM have high-dimensional token representations. In high-dimensional spaces, nearest neighbor search often suffer from the curse of dimensionality. To reduce this problem, we apply dimension reduction to NN-UE similar to He et al. (2021a). In this study, we use Principal Component Analysis (PCA) as a dimension reduction algorithm to reduce the dimension of the datastore representations and the query representation .
Results: Combination of PQ, Clustering, and Dimension Reduction
We evaluate the uncertainty estimation performance and inference speed when applying PQ, clustering, and dimension reduction are applied sequentially. The evaluations are performed on the OntoNotes 5.0 test set, and the results for different parameters of PQ, clustering and dimension reduction are shown in Appendix C. Table 8 shows the results on OntoNotes 5.0 and as ID/OOD, respectively. We can see that while the uncertainty performance is not significantly degraded when PQ and clustering are applied simultaneously to NN-UE, ECE and MCE are degraded when dimension reduction by PCA is further applied.666Distance recomputation does not mitigate this behavior, see Appendix D. On the other hand, E-AURC is not degraded when applying the combination of PQ, clustering, and dimension reduction. Also, we can obtain the improvement of inference speed rivaling the level of speed close to SR by applying PCA.
To deepen our understanding of the changes in the behavior of the uncertainty performance due to appling of approximate NN search techniques or dimension reduction in NN-UE, we calculated the coverage that how much the indices obtained when using the default exhaustive search are covered when applying PQ, clustering, and dimension reduction, sequentially. Table 9 shows the coverages on OntoNotes 5.0 and as ID/OOD settings, respectively.
We can see that while applying PQ and clustering simultaneously or applying PCA individually covers a certain percentage of the example indices obtained from NN-UE, which uses exhaustive search by default, applying PQ, clustering, and PCA simultaneously hardly covers any of the indices from the default NN-UE. It is assumed that applying PQ and PCA leads to coarse distance computation in a single subvector, which would correspondingly degrade the uncertainty estimation performance in NN-UE. On the other hand, if they are not applied simultaneously, the degradation in the uncertainty performance is limited.
8 Conclusion
In this paper, we proposed NN-UE, which estimates uncertainty by using the distance to neighbors and labels of neighbors. The experimental results showed that our method showed higher uncertainty estimation performance than existing uncertainty estimation methods in SA, NLI and NER. Furthermore, we showed that the uncertainty estimation performance can be further improved by adjusting the number of neighbors in NN-UE. Our method can greatly improve uncertainty estimation performance, especially in text classification tasks, with little degrading in inference speed. On the other hand, to address the degradation of the inference speed in token-level tasks such as NER, we investigated the effects of efficient neighbor search techniques in NN-UE. As a result, we found that product quantization, clustering, or dimension reduction improves inference speed without degrading the uncertainty estimation much more, unless combining all of them simultaneously.
9 Limitations
In this study, we focused only on the classification-based tasks. On the other hand, taking advantage of the recent growth of Large Language Models, uncertainty estimation in text generation is also attracting attention Fadeeva et al. (2023); Lin et al. (2024). Therefore, to investigate the effectiveness of NN-UE in text generation tasks is an interesting direction for future research. Furthermore, although NN-UE only used the representation of the last layer of the base model, exploring for an appropriate representation for uncertainty estimation is a future challenge.
Ethical Considerations
In this study, we used existing datasets that have cleared ethical issues following policies of published conferences. Therefore, they do not introduce any ethical problems. On the other hand, we have an ethical consideration about uncertainty estimation. Specifically, decision support systems with machine learning algorithms do not necessarily have a positive effect on performance. Jacobs et al. (2021) showed that collaboration with machine learning models does not significantly improve clinician’s treatment selection performance, and that performance is significantly degraded due to the presentation of incorrect recommendations. This problem is expected to remain even if uncertainty estimation methods are applied to machine learning models. In addition, introducing uncertainty estimation methods could conversely lead humans to give overconfidence in machine learning models, resulting in performance degradation.
Acknowledgements
The authors also acknowledge the Nara Institute of Science and Technology’s HPC resources made available for conducting the research reported in this paper.
Appendix A Training Settings for Density Estimator in Density Softmax
In Density Softmax Bui and Liu (2024a), we use RealNVP Dinh et al. (2017) which has two coupling structures. Table 10 shows the hyperparameters for training RealNVP as the density estimator in Density Softmax.
| Hyperparameters | Values |
| learning rate | 1e-4 |
| optimizer | AdamW Loshchilov and Hutter (2019) |
| early stopping patient | 5 |
| number of coupling layers | 4 |
| hidden units | 16 |
Appendix B Details of Evaluation Metrics
Expected Calibration Error (ECE)
ECE Naeini et al. (2015) quantifies the difference between the accuracy and confidence of a model. Formally, ECE is expressed as:
where is the number of confidence interval bins, denotes the set of examples with predicted confidence scores in the -th bin, is the total number of examples, is the accuracy of the model on the examples in , and is the average confidence of the model on the examples in . In this study, we use .
Maximum Calibration Error (MCE)
MCE, as detailed by Naeini et al. (2015) measures the maximum difference between the model’s accuracy and the confidence across variouts confidence levels. MCE is defined as:
A lower MCE means that there is a small risk that the confidence of the model’s prediction will deviate greatly from the actual correct answer. In this study, we use , same as ECE.
Area Under the Risk-Coverage curve (AURC)
The AURC is the area of the risk-coverage curve when the confidence levels of the forecasts corresponding to the data points are sorted in descending order. The larger the area, the lower the error rate corresponding to a higher confidence level, which means that the output confidence level is more appropriate. Formally, AURC is defined as:
where returns 1 if the prediction is wrong and 0 otherwise.
Excess-Area Under the Risk-Coverage curve (E-AURC)
E-AURC Geifman et al. (2019) is a measure of the AURC score normalized by the smallest risk-coverage curve area , where is the error rate of the model. The reason for normalizing the AURC is that the AURC depends on the predictive performance of the model and allows for performance comparisons of confidence across different models and training methods. E-AURC is defined as:
| E-AURC | (9) |
E-AURC scores are reported with multiplying by 1,000 due to visibility.
Appendix C Each Result of Product Quantization, Clustering, and Dimension Reduction
C.1 Product Quantization
We evaluated uncertainty estimation performance and inference time when the number of clusters in the codebook was fixed at 32, and the number of subvectors was changed to .
Table 11 shows the uncertainty estimation performance and inference time results in different . In ECE and E-AURC, there are almost no degradation in uncertainty estimation performance due to PQ. On the other hand, in MCE in ID setting, the uncertainty estimation performance consistently degrades. Furthermore, compared to NN-UE among different , the larger , the better the uncertainty estimation performance tends to improve, but the inference time increases.
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| OntoNotes 5.0 (In-domain) | ||||
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 2.490.08 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 4.940.10 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 |
| NN-UE () | 1.900.27 | 31.1811.17 | 20.161.12 | 3.270.06 |
| NN-UE () | 1.960.31 | 31.3318.74 | 20.231.27 | 3.320.05 |
| NN-UE () | 1.880.34 | 31.0616.36 | 20.161.23 | 4.110.11 |
| OntoNotes 5.0 (Out-of-domain) | ||||
| SR | 17.050.69 | 37.063.13 | 81.494.17 | 5.750.27 |
| NN-UE (w/o label) | 8.780.62 | 24.911.81 | 70.104.03 | 10.360.21 |
| NN-UE | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| NN-UE () | 7.660.48 | 17.073.81 | 74.475.53 | 7.220.19 |
| NN-UE () | 7.570.45 | 16.432.73 | 74.385.36 | 7.230.16 |
| NN-UE () | 7.570.44 | 16.382.66 | 74.355.49 | 8.900.18 |
C.2 Clustering
In this study, we evaluate the uncertainty estimation performance and inference speed when the number of clusters is fixed and number of cluster centroids to search changes.
Table 12 shows the performance of uncertainty estimation when changing in ID and OOD settings using OntoNotes 5.0. In ECE, scores are slightly reduced for ID, but only slightly worse for OOD; MCE also shows degradation for ID but little for OOD, and even improves when ; E-AURC shows almost no change in scores when is changed for both ID and OOD. In terms of inference time, the larger , the longer it takes.
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| OntoNotes 5.0 (In-domain) | ||||
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 2.490.08 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 4.940.10 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 |
| NN-UE () | 1.820.28 | 30.1816.77 | 20.141.21 | 2.840.08 |
| NN-UE () | 1.860.25 | 29.4816.91 | 20.131.21 | 3.110.03 |
| NN-UE () | 1.920.31 | 28.5511.24 | 20.131.22 | 3.310.06 |
| NN-UE () | 1.830.28 | 27.009.43 | 20.141.21 | 3.710.06 |
| OntoNotes 5.0 (Out-of-domain) | ||||
| SR | 17.050.69 | 37.063.13 | 81.494.17 | 5.750.27 |
| NN-UE (w/o label) | 8.780.62 | 24.911.81 | 70.104.03 | 10.360.21 |
| NN-UE | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| NN-UE () | 7.520.41 | 16.011.92 | 74.335.37 | 6.090.28 |
| NN-UE () | 7.560.36 | 16.933.38 | 74.315.39 | 6.650.17 |
| NN-UE () | 7.600.41 | 17.122.35 | 74.345.35 | 7.330.21 |
| NN-UE () | 7.530.40 | 17.282.45 | 74.335.37 | 7.890.12 |
C.3 Dimension Reduction
As shown in Table 13, the uncertainty estimation performance depends on the number of target dimension, and the performance degrades when or . On the other hand, the performance in is almost the same as default NN-UE. This is because while dimension reduction mitigates the problem of high dimensionality, the failure to extract adjacent examples by NN search when applying PCA is less likely compared to PQ and clustering if the target dimension in dimension reduction is not reduced too much.
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| OntoNotes 5.0 (In-domain) | ||||
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 2.490.08 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 4.940.10 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 |
| NN-UE () | 1.890.37 | 31.0114.35 | 20.061.25 | 3.240.08 |
| NN-UE () | 1.800.36 | 27.8513.80 | 20.131.29 | 3.410.10 |
| NN-UE () | 1.800.40 | 26.2312.61 | 20.131.28 | 3.850.06 |
| OntoNotes 5.0 (Out-of-domain) | ||||
| SR | 17.050.69 | 37.063.13 | 81.494.17 | 5.750.27 |
| NN-UE (w/o label) | 8.780.62 | 24.911.81 | 70.104.03 | 10.360.21 |
| NN-UE | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| NN-UE () | 7.480.41 | 16.202.75 | 74.335.49 | 7.370.26 |
| NN-UE () | 7.540.45 | 16.422.73 | 74.305.44 | 7.750.24 |
| NN-UE () | 7.560.43 | 16.132.59 | 74.265.40 | 8.510.46 |
Appendix D Distance Recomputation for NN-UE
When using efficient NN search techniques in Section 7.2, we use approximate distances to compute Eq. 5. Although we can get raw vectors by using the example indices obtained from approximate nearest neighbor search and compute accurate distance, in NN-LM this has been shown to lead to performance gains and latency degradation He et al. (2021a). We measure the uncertainty estimation performance and inference speed when PQ, clustering, and dimension reduction are applied simultaneously and re-computing accurate distances, reported in Table 14. These results show that the uncertainty estimation performance does not improve except for MCE in the ID setting, and the latency is about 5-7x slower when reading raw vectors from the datastore and re-computing distances.
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| OntoNotes 5.0 (In-domain) | ||||
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 |
| NN-UE (Approx.) | 2.140.37 | 33.5210.84 | 20.121.26 | 2.870.04 |
| NN-UE (Recomp.) | 2.350.44 | 30.477.50 | 20.161.17 | 16.240.77 |
| OntoNotes 5.0 (Out-of-domain) | ||||
| NN-UE | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| NN-UE (Approx.) | 8.080.53 | 24.035.46 | 74.505.42 | 6.200.20 |
| NN-UE (Recomp.) | 8.300.51 | 25.675.26 | 74.585.53 | 34.220.78 |
Appendix E Licenses of Datasets, Tools and Models
Datasets
IMDb movie dataset can be used for research purpose as described in https://developer.imdb.com/non-commercial-datasets/. Yelp Polarity dataset can be used for academic purpose as described in https://s3-media0.fl.yelpcdn.com/assets/srv0/engineering_pages/f64cb2d3efcc/assets/vendor/Dataset_User_Agreement.pdf. MNLI dataset is licensed for research purpose as described in Williams et al. (2018). SNLI dataset can be used for research purpose as described in https://nlp.stanford.edu/projects/snli/. OntoNotes 5.0 dataset can be used for research purpose as described in https://catalog.ldc.upenn.edu/LDC2013T19.
Tools
is licensed by Apache-2.0. is MIT-licensed.
Models
and from Huggingface model checkpoints are MIT-licensed.