### v2

# Efficient Nearest Neighbor based Uncertainty Estimation for Natural Language Processing Tasks

Abstract
Trustworthiness in model predictions is crucial for safety-critical applications in the real world. However, deep neural networks often suffer from the issues of uncertainty estimation, such as miscalibration. In this study, we propose -Nearest Neighbor Uncertainty Estimation (NN-UE), which is a new uncertainty estimation method that uses not only the distances from the neighbors, but also the ratio of labels in the neighbors. Experiments on sentiment analysis, natural language inference, and named entity recognition show that our proposed method outperforms the baselines and recent density-based methods in several calibration and uncertainty metrics. Moreover, our analyses indicate that approximate nearest neighbor search techniques reduce the inference overhead without significantly degrading the uncertainty estimation performance when they are appropriately combined.
Efficient Nearest Neighbor based Uncertainty Estimation for Natural Language Processing Tasks
Wataru Hashimoto, Hidetaka Kamigaito, Taro Watanabe Nara Institute of Science and Technology {hashimoto.wataru.hq3, kamigaito.h, taro}@is.naist.jp
1 Introduction
In order to deploy Deep Neural Networks (DNNs) including Pre-trained Language Models (PLMs) in safety-critical areas, uncertainty estimation (UE) is important. Improving the predictive uncertainty will calibrate the prediction Guo et al. (2017),111”Calibration” means the confidence of the prediction aligns with its accuracy. or enhance the selective prediction performance which reduces incorrect predictions by providing the option to abstain from the model prediction Galil et al. (2023). On the other hand, DNNs often fail to quantify the predictive uncertainty, for example, causing miscalibrated prediction Guo et al. (2017). Such UE performance problems can be mitigated by the PLMs, such as BERT Devlin et al. (2019) or DeBERTa He et al. (2021b), that are self-trained on vast amounts of data Ulmer et al. (2022); nevertheless, there remains considerable room for improvement Desai and Durrett (2020).
To address the challenge of UE, multiple stochastic inferences such as MC Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017) are generally effective. On the other hand, these methods require multiple stochastic inferences for a single data instance, which leads to high computational cost, and makes them impractical for real world application. To balance reasonable predictive uncertainty with computational efficiency, Temperature Scaling Guo et al. (2017) , which scales logits by a temperature parameter, is commonly employed. Furthermore, density-based methods, such as Density Softmax Bui and Liu (2024) and Density Aware Calibration (DAC) Tomani et al. (2023), have demonstrated promising UE performance and inference costs by adjusting model outputs based on estimated density.
However, both Density Softmax and DAC only use the density of the training data. Relying on density alone can sometimes lead to overconfident predictions, even when such confidence is unwarranted. For instance, the neighbor of the input may contain many examples with labels that differ from the predicted label. In this situation, the prediction should obviously not be trusted. Therefore, we hypothesized that considering both the density and the label information of the neighbors will improve UE performance.
In this study, we propose -Nearest Neighbor Uncertainty Estimation (NN-UE), a new density-based UE method that reflects nearest neighbor labels. As illustrated in Figure 1, NN-UE is designed to achieve the highest prediction confidence when the input and the nearest neighbors are both close in distance and share the same label as the predicted label. Our method weights logits according to the score from the distance between the input example and its neighbors in the datastore created by the training data and the ratio of the model’s predicted label matched with the labels in the neighbors. In addition, our method requires only a single forward inference of the model with almost no additional computational cost. The contributions of this research are as follows.
First, our experiments show that NN-UE improves the UE performance of existing baselines in sentiment analysis, natural language inference, and named entity recognition in both in-domain and out-of-domain settings by combining neighbor label information and distances from neighbors. On the other hand, we also find that naive NN-UE makes less efficient for token-level tasks such as sequence-labeling based named entity recognition due to the execution of NN to each token.
Second, to mitigate the above latency problem in NN-UE, we show that approximate NN search or dimension reduction in NN-UE improves the inference speed without degrading UE performance much more, while combining them leads to degrading the UE performance.
Our code is available at https://github.com/wataruhashimoto52/knn_ue.
2 Related Work
Uncertainty Estimation for Natural Language Processing Tasks
Studies about UE for NLP tasks are limited when compared with those for image datasets. Kotelevskii et al. (2022) has shown excellent performance in classification with rejection tasks and out-of-distribution detection tasks using uncertainty scores using density estimation results. Vazhentsev et al. (2022) performed misclassification detection using Determinantal point processes Kulesza and Taskar (2012), spectral normalization, Mahalanobis distance and loss regularization in text classification and NER. However, these are still focusing only on the feature representation or the density, not the labels of the neighbors. He et al. (2024) proposed a framework that considers uncertainty between tokens in NER. However, the target task is limited to NER, and it is not for confidence calibration. Hashimoto et al. (2024) shows that simple data augmentation methods in NER can improve UE performance without additional inference costs, but its effectiveness is limited in the in-domain. Our NN-UE improves the UE performance in the in-domain and the out-of-domain classification and NER tasks using not only NN density but also neighbor labels.
-Nearest Neighbor Language Models / Machine Translation
-Nearest Neighbor Language Model (NN-LM) Khandelwal et al. (2020) has been proposed, which performs linear interpolation of NN probability based on distance from neighbors and base model probability, in the language modeling task. -Nearest Neighbor Machine Translation (NN-MT) applied the NN-LM framework to machine translation Khandelwal et al. (2021). NN-LM and NN-MT have been successful because they enhance predictive performance through the memorization and use of rich token representations of pre-trained language models and mitigate problems such as a sparsity comes from low-frequency tokens Zhu et al. (2023). The main issue on NN-LM and NN-MT is the inference overhead, and there are several studies to solve this problem. He et al. (2021a) employs datastore compression, adaptive retrieval, and dimension reduction to reduce computational overhead with retaining perplexity. Deguchi et al. (2023) dramatically improves decoding speed by dynamically narrowing down the search area based on the source sentence. We investigate that whether UE performance in NN-UE can keep or not with reducing inference time by introducing some of the speed-up techniques established in kNN-LM/MT.
3 Preliminary
3.1 Definitions
In multiclass classification, we assume a dataset consisting of examples, where denotes its corresponding class label among possible classes.222In the case of sequence labeling, we can interpret the number of data as the product of the raw number of data instances and the sequence length. We use the trained neural network feature extractor and the classifier for classification, where . gives us the logits and we obtain the confidence .
3.2 Density Softmax
Density Softmax Bui and Liu (2024) obtains confidence by weighting logits with normalized log-likelihood from a trained density estimator. are the parameters of the density estimator; is the normalized log-likelihood from the density estimator, then the corrected confidence is written as
In Density Softmax, the closer the normalized log-likelihood to zero, the closer the prediction to Uniform distribution. Density Softmax achieves reasonable latency and competitive UE performance with state-of-the-art methods at the cost of demanding the density estimator training and multiple base model training.333Details for the density estimator in this study are in Appendix B.
3.3 Density Aware Calibration (DAC)
DAC is a confidence calibration method using multiple feature representations, which is similar to the NN-based out-of-distribution detection Sun et al. (2022). DAC Tomani et al. (2023) scales the logits by using sample-dependent temperature
| where | ||||
are the weights for every layer of the base model, is the averaged distance from NN search on -th layer, and is the bias term. are optimized using the L-BFGS-B method Liu and Nocedal (1989) based on the loss in the validation set. In the original DAC paper, the UE performance tends to improve with the increase in the number of layer representation Tomani et al. (2023). Therefore, we use all the hidden representations in each layer of the base PLMs.
DAC is a non-parametric method that makes not assumptions about the training data distribution unlike Density Softmax Bui and Liu (2024), which relies on some density estimators. On the other hand, the recent NN-based DAC still relies only on the distances to the neighbors. These methods do not take into account the label information of the input neighbors, which limits the improvement of the UE performance.
4 Proposed Method: -Nearest Neighbor Uncertainty Estimation (NN-UE)
The main idea of our proposed method, NN-UE, stems from the notion that the density-based UE methods can be further improved by using label information about the training data instances that make up the density.
In order to take into account the variance of neighbor labels, our NN-UE explicitly includes the label agreement information of the predicted instance and its neighbor examples when calculating the confidence. More specifically, we regard the prediction as more reliable only when the prediction is in a region where training data is dense and the predicted label and the labels of the data instances that make up the dense region are mostly the same, as illustrated in the right part of Figure 1. Otherwise, for example, if there is a lot of discrepancy in the neighbor labels and the predicted label, we treat the prediction as unreliable, indicated in the middle of Figure 1.
In our NN-UE, we introduce two terms: one related to the density of the training data and one related to the degree of agreement of the predicted data and neighbor labels. Confidence of -th label obtained by NN-UE is following the formula:
| where | ||||
is the number of neighbors from NN search, is the count when the predicted label and the label of the -th neighbor is same, is the distance between the -th representation obtained by NN search and the representations of training data.444Note that NN-UE is also ”accuracy-preserving” same as DAC because is a scalar, not a class-wise score. The parameters , , and are optimized using the L-BFGS-B method based on the loss in the validation set.
When the distance and label terms are smaller and is closer to zero, the closer the prediction is to Uniform distribution, which allows us to better estimate the confidence of the prediction. In this study, we also conduct experiments without the label term in Equation 5, to emphasize the importance of NN neighbor labels in UE. We summarize a diagram of NN-UE in Figure 2.
5 Experimental Settings
5.1 Tasks and Datasets
We measure the UE performance on Sentiment Analysis (SA), Natural Language Inference (NLI), and Named Entity Recognition (NER) in In-domain (ID) and Out-of-Domain (OOD) settings.555The datasets in SA and NLI were set up with reference to Xiao et al. (2022). Dataset statistics are described in Appendix A.
Sentiment Analysis (SA)
Natural Language Inference (NLI)
Named Entity Recognition (NER)
extracts the named entities, such as a person, organization, or location. The NER task was carried out in the framework of sequence labeling. We regard the OntoNotes 5.0 dataset Pradhan et al. (2013) broadcast news () domain as ID, and newswire () and telephone conversation () domains as OOD.
5.2 Existing Methods
We employ the simple baselines: Softmax Response (SR) Cordella et al. (1995), Temperature Scaling (TS) Guo et al. (2017), Label Smoothing Miller et al. (1996); Pereyra et al. (2017) and MC Dropout Gal and Ghahramani (2016). In addition, we use the recent strong baselines for UE: Spectral-Normalized Gaussian Process (SNGP) Liu et al. (2020), Posterior Networks (PN) Charpentier et al. (2020), Mahalanobis Distance with Spectral-Normalized Network (MDSN) Vazhentsev et al. (2022), E-NER Zhang et al. (2023), Density Softmax Bui and Liu (2024), and DAC Tomani et al. (2023). Details on baselines can be found in Appendix C. We have also experimented with a variant of NN-UE without the label term in Eq. 5, denoted by “w/o label” to emphasize the impact of the neighbor labels.
5.3 Training Settings
In all experiments, we train and evaluate the models on a single NVIDIA A100 GPU with 40GB of memory. We used 666 https://huggingface.co/microsoft/deberta-v3-base and 777https://huggingface.co/microsoft/mdeberta-v3-base He et al. (2023), as the Transformer encoder from Wolf et al. (2020) pre-trained model checkpoints. We use the cross-entropy loss in all experiments, including the optimization of hyperparameters in NN-UE. Batch size is 32, and the initial learning rate was set to 1e-5. The gradient clipping is applied with the maximum norm of 1. All experiments are run five times, and we report the mean and standard deviation of the scores.
Datastore Construction
It is necessary to maintain the representation of the data for training a density estimator in Density Softmax and NN search in DAC and NN-UE. We use the final layer representations corresponding to CLS tokens in SA and NLI. In NER, we stored the hidden representation of the final layer as a token representation corresponding to the beginning of the word.
-Nearest Neighbor Search
We use Douze et al. (2024) as the GPU-accelerated NN search toolkit. Unless otherwise specified, we fix the number of neighbors in NN search,888In Section 7.1, we conducted experiments to examine the behavior when varying over the set , with representing the median. and use which is an index for exact search in L2 norm, as the default in NN-UE.
| Methods | IMDb (In-domain) | Yelp (Out-of-domain) | ||||||
| ECE (↓) | MCE (↓) | AUROC (↑) | E-AURC (↓) | ECE (↓) | MCE (↓) | AUROC (↑) | E-AURC (↓) | |
| SR | 4.420.41 | 24.063.52 | 98.350.10 | 10.602.81 | 4.691.20 | 21.026.74 | 98.150.39 | 11.843.15 |
| TS | 4.100.31 | 20.435.01 | 98.450.21 | 11.362.82 | 5.101.19 | 19.701.35 | 98.200.46 | 12.914.12 |
| LS | 1.880.41 | 21.504.53 | 98.360.45 | 14.527.24 | 2.530.43 | 16.473.51 | 98.300.45 | 12.906.09 |
| MC Dropout | 4.280.27 | 23.743.52 | 98.570.12 | 9.171.74 | 4.330.54 | 20.172.79 | 98.280.25 | 10.012.01 |
| SNGP | 4.180.30 | 22.694.83 | 98.530.15 | 9.951.17 | 4.890.59 | 21.284.68 | 98.100.27 | 11.422.14 |
| PN | 4.280.43 | 24.430.20 | 98.060.27 | 10.995.63 | 4.690.35 | 24.410.32 | 97.560.25 | 15.823.94 |
| MDSN | 4.450.43 | 23.975.05 | 98.480.08 | 10.250.86 | 5.320.92 | 21.332.91 | 98.000.20 | 11.123.53 |
| Density Softmax | 4.230.36 | 27.106.92 | 98.340.08 | 11.392.48 | 4.990.48 | 21.983.68 | 98.090.24 | 13.052.72 |
| DAC | 1.510.33 | 14.172.73 | 98.360.37 | 12.726.15 | 2.350.12 | 6.442.23 | 97.860.60 | 14.265.90 |
| NN-UE (w/o label) | 1.330.36 | 13.133.24 | 98.650.13 | 9.360.36 | 2.230.29 | 6.332.76 | 98.270.11 | 10.970.91 |
| NN-UE | 0.950.12† | 9.021.39† | 98.640.12 | 7.970.61† | 1.450.15† | 4.171.52 | 98.230.39 | 9.920.61 |
| Methods | MNLI (In-domain) | SNLI (Out-of-domain) | ||||||
| ECE (↓) | MCE (↓) | AUROC (↑) | E-AURC (↓) | ECE (↓) | MCE (↓) | AUROC (↑) | E-AURC (↓) | |
| SR | 8.360.61 | 37.617.53 | 97.030.12 | 31.292.23 | 9.770.55 | 36.6114.05 | 96.070.17 | 37.620.67 |
| TS | 2.731.86 | 15.8111.05 | 97.060.02 | 31.241.86 | 3.921.79 | 18.1310.69 | 96.080.13 | 38.402.06 |
| LS | 2.890.14 | 28.647.90 | 96.560.55 | 37.9812.64 | 3.970.45 | 23.186.17 | 95.610.40 | 44.189.18 |
| MC Dropout | 8.130.65 | 30.176.83 | 96.970.06 | 32.312.25 | 9.620.53 | 28.905.03 | 96.100.11 | 37.192.99 |
| SNGP | 10.450.56 | 35.4213.89 | 95.910.12 | 42.032.72 | 14.281.04 | 31.163.42 | 93.400.44 | 63.216.84 |
| PN | 33.830.51 | 37.100.71 | 96.960.10 | 26.331.22 | 32.010.61 | 35.370.58 | 95.570.29 | 40.944.49 |
| MDSN | 8.340.46 | 29.046.43 | 97.070.14 | 32.032.29 | 9.440.47 | 38.5913.94 | 96.110.12 | 38.913.06 |
| Density Softmax | 8.420.43 | 36.205.78 | 97.030.10 | 32.563.29 | 10.090.40 | 33.594.57 | 95.960.19 | 41.432.25 |
| DAC | 1.420.30 | 18.7910.81 | 96.920.10 | 33.892.60 | 2.270.16 | 11.553.48 | 96.080.07 | 40.233.00 |
| NN-UE (w/o label) | 1.280.43 | 16.5311.45 | 97.090.10 | 30.222.80 | 2.120.36 | 10.006.07 | 96.120.16 | 37.334.70 |
| NN-UE | 1.410.47 | 10.772.34† | 97.180.09 | 23.831.29† | 1.800.37 | 5.121.47† | 96.000.22 | 34.972.48 |
5.4 Evaluation
To evaluate the confidence calibration performance, we choose Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) Naeini et al. (2015). For selective prediction, we evaluate Area Under the Receiver Operator Characteristic curve (AUROC) and Excess-Area Under the Risk-Coverage curve (E-AURC) Geifman et al. (2019). Evaluation metrics computation details are described in Appendix D. In NER, we performed the evaluation of the UE performance with the flat recombination of the labels and the confidence for all tokens, respectively.
6 Results
6.1 Sentiment Analysis
In SA, we evaluate the confidence calibration, selective prediction and out-of-distribution detection performance.
Confidence Calibration and Selective Prediction
First, we present the UE results for sentiment analysis by differentiating the in-domain and out-of-main performance in Table 1. NN-UE consistently outperforms existing methods in terms of ECE, MCE, and E-AURC. In AUROC, LS outperforms in OOD setting, but NN-UE outperforms existing methods in ID setting. Furthermore, the proposed method clearly outperforms DAC that uses neighbor search results for each hidden representation with the additional label term. The lower UE performance than NN-UE in DAC is probably due to the difficulty in optimizing hyperparameters by comprising many layers.
Out-of-Distribution Detection
Following the previous study Tomani et al. (2023), we carried out the experiments in the out-of-distribution detection task, which determines whether a data instance is in-domain or not. This task is based on the intuition that we want to return predictions with high confidence in ID but with low confidence in predictions in OOD. We evaluated the out-of-distribution detection performance by using maximum softmax probability as the uncertainty score, and report FPR@95 (the FPR when the TPR is 95%), AUROC, Area Under the Precision-Recall curve (AUPR)-in and AUPR-out. AUPR-in indicates the AUPR score when ID samples are treated as positive; AUPR-out is vice versa.
Table 3 shows the out-of-distribution detection results when using IMDb/Yelp datasets as ID/OOD, respectively, in model. NN-UE consistently shows the out-of-distribution detection performance improvement.
| Methods | FPR@95 (↓) | AUROC (↑) | AUPR-In (↑) | AUPR-Out (↑) |
| SR | 82.519.49 | 63.185.14 | 69.512.57 | 54.708.48 |
| TS | 83.127.50 | 65.633.64 | 70.992.02 | 56.196.11 |
| LS | 86.884.27 | 62.172.83 | 69.501.51 | 51.383.81 |
| MC Dropout | 87.333.38 | 63.964.09 | 70.132.39 | 53.185.41 |
| SNGP | 81.923.46 | 63.273.07 | 68.832.10 | 55.913.20 |
| PN | 82.845.11 | 67.544.29 | 66.592.45 | 55.325.26 |
| Density Softmax | 87.543.14 | 58.734.33 | 67.342.57 | 49.194.36 |
| DAC | 84.984.19 | 64.656.18 | 70.693.59 | 54.817.29 |
| NN-UE (w/o label) | 75.872.16 | 70.441.70 | 74.771.44† | 63.392.24 |
| NN-UE | 73.555.01† | 71.112.92† | 73.802.19 | 65.013.45† |
6.2 Natural Language Inference
We show the results of in-domain and out-of-domain UE in NLI task using the DeBERTaV3 model in Table 2. Similar to Section 6.1, NN-UE shows the best UE performance, especially when the label term is included. Galil et al. (2023) have reported that improving calibration performance does not necessarily lead to the improved selective prediction performance, but our proposed method improves both types of metrics. On the other hand, the degree of improvement is larger for calibration performance. Specifically, the largest improvement is obtained on SNLI, where NN-UE reduces MCE by more than 31.49 % compared to SR. Additional experimental results on the Brier score can be found in Appendix E.
| Methods | (In-domain) | (Out-of-domain) | (Out-of-domain) | ||||||
| ECE (↓) | MCE (↓) | E-AURC (↓) | ECE (↓) | MCE (↓) | E-AURC (↓) | ECE (↓) | MCE (↓) | E-AURC (↓) | |
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 17.050.69 | 37.063.13 | 81.494.17 | 21.202.03 | 42.605.84 | 76.055.72 |
| TS | 5.340.43 | 75.7121.96 | 19.631.22 | 12.760.62 | 26.573.97 | 72.904.72 | 19.690.95 | 47.727.34 | 71.878.83 |
| LS | 6.460.74 | 50.9926.73 | 24.931.19 | 14.780.61 | 30.542.84 | 81.506.98 | 20.992.16 | 65.4017.16 | 76.657.33 |
| MC Dropout | 6.760.64 | 53.1326.07 | 19.913.39 | 15.271.01 | 33.604.93 | 77.213.72 | 21.931.63 | 56.5612.32 | 75.689.30 |
| E-NER | 7.980.42 | 61.8727.06 | 19.441.81 | 17.420.88 | 40.465.33 | 74.324.47 | 25.422.09 | 59.1610.33 | 72.006.57 |
| Density Softmax | 7.320.25 | 59.0527.76 | 25.172.63 | 16.100.62 | 44.6621.67 | 80.148.50 | 24.401.84 | 62.5010.46 | 80.066.27 |
| DAC | 1.620.42 | 42.9628.25 | 21.472.90 | 7.910.75 | 25.285.15 | 75.242.43 | 14.421.57 | 47.9220.98 | 80.728.19 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66† | 8.780.62 | 24.911.81 | 70.104.03 | 14.610.67 | 35.267.16† | 65.418.11 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 7.500.42 | 16.532.61† | 74.275.43 | 14.150.33 | 39.846.02 | 71.819.04 |
6.3 Named Entity Recognition
To evaluate NLP tasks other than simple multi-class classification, we evaluate NN-UE in NER. Since NER focuses on entities, we use the product of the confidence of the tokens that construct a single entity as the confidence of the entity.
Table 4 shows the results of in-domain and out-of-domain UE using the OntoNote 5.0 dataset in . NN-UE shows the best performance in 4 cases, i.e., ECE or MCE, often resulting in large improvements over SR. On the other hand, E-AURC in NER is consistently better without using the NN-UE label term.999Label imbalance or large number of class can significantly affect E-AURC on NER when using kNN-UE with the label term. Details are in Appendix F. E-NER, a recent UE method specifically designed for NER, is close to NN-UE in its entity level selective prediction performance, but the calibration performance is not high.
NN-UE shows good UE performance even when the target domain is relatively far from source domain , such as . We have hypothesized that NN-UE might not work if the prediction target is too far from the training data distribution. If the prediction target is too far from the training data, the representation of the prediction from the model will be unreliable when compared to the prediction in the same domain as the training data. In general, methods based on feature distances assume that they maintain information relevant to the correctness of the prediction Postels et al. (2022). Our experiments have shown that the problem could be mitigated probably because the domains that the base models do not recognize are limited in the NLP community where there are many strong pre-trained models based on self-supervised learning such as DeBERTaV3.
| Scores | Correct Instances | Incorrect Instances | ||
| NN-UE (w/o label) | NN-UE | NN-UE (w/o label) | NN-UE | |
| 0.49 | 0.50 | 0.41 | 0.27 | |
| Confidence | 0.95 | 0.93 | 0.82 | 0.72 |
| Methods | MNLI | OntoNotes 5.0 |
| SR | 8.410.03 | 2.490.08 |
| TS | 8.420.07 | 2.510.08 |
| LS | 8.440.06 | 2.530.03 |
| MC Dropout | 157.520.51 | 39.810.39 |
| SNGP | 10.582.09 | - |
| PN | 9.110.07 | - |
| MDSN | 9.651.36 | - |
| E-NER | - | 2.510.12 |
| Density Softmax | 8.570.06 | 2.590.05 |
| DAC | 785.156.72 | 183.460.76 |
| NN-UE (w/o label) | 9.050.07 | 4.940.10 |
| NN-UE | 9.080.10 | 4.990.07 |
7 Analysis
7.1 Impact of Top-
To understand the behavior of NN-UE, we evaluated the performance in UE when changing the number of neighbors during NN execution.
Figure 3 and 4 show the results for SA and NER, respectively. As is noticeable in NER, the smaller , the better UE tends to be. These results suggest that our method requires that nearer examples to calibrate confidence, but more distant examples are not important. When calculating in Eq. 5, automatically adjusting the importance weights based on the order or distance of the retrieved nearest neighbors could further improve UE performance. Similar experimental and theoretical analysis of out-of-distribution detection using only NN distance also suggests that using -th example is preferable Sun et al. (2022). Providing a similar theoretical justification for our NN-UE is an interesting future direction.
| OntoNotes 5.0 (In-domain) | OntoNotes 5.0 (Out-of-domain) | |||||||
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 2.490.08 | 17.050.69 | 37.063.13 | 81.494.17 | 5.750.27 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 4.940.10 | 8.780.62 | 24.911.81 | 70.104.03 | 10.360.21 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| + PQ | 1.960.31 | 31.3318.74 | 20.231.27 | 3.320.05 | 7.570.45 | 16.432.73 | 74.385.36 | 7.230.16 |
| + IVF | 1.920.31 | 28.5511.24 | 20.131.22 | 3.310.06 | 7.600.41 | 17.122.35 | 74.345.35 | 7.330.21 |
| + DR | 2.140.37 | 33.5210.84 | 20.121.26 | 2.870.04 | 8.080.53 | 24.035.46 | 74.505.42 | 6.200.20 |
| Methods | OntoNotes 5.0 | OntoNotes 5.0 |
| NN-UE | 100.0 | 100.0 |
| + PQ | 21.30 | 51.68 |
| + IVF | 18.60 | 11.04 |
| + DR | 0.02 | 0.04 |
| Only DR | 43.98 | 20.35 |
7.2 Importance of Label Term in
We analyze the impact of the label term Eq. 5 on the NN-UE confidence computation. We have shown that the UE performance is improved in several experiments. However, it is not obvious whether the improvement in UE performance is due to the reduction in primarily caused by adding the label term. Therefore, we examined the values for correctly and incorrectly predicted instances in both the absence and presence of the label term in NN-UE. Table 5 shows the distance terms, label terms and results for each case. If the predictions are correct, the growth of due to the label term is limited. On the other hand, NN-UE with label term remarkably reduce leading to the reduced confidence when the predictions are incorrect. This result suggests that the improvement of the evaluation metrics in NN-UE with label term is not achieved by increasing the confidence when the prediction is correct, but by appropriately reducing the confidence when the prediction is incorrect.
7.3 Impact of Efficient Nearest Neighbor Search Techniques
We investigate the inference time and UE performance when applying approximate nearest neighbor search techniques and dimension reduction when executing NN search in NN-UE as a real world application. As shown in Table 6,101010Other results can be found in Appendix G. in the sequence labeling based NER, which requires executing NN searches per token, it takes twice as much inference time as SR.111111Inference times do not increase as dramatically as -Nearest Neighbor Language Model Khandelwal et al. (2020) because NN can be executed in parallel for both classification and NER. On the other hand, in -Nearest Neighbor Language Model (NN-LM) Khandelwal et al. (2020), dimension reduction and approximate NN search techniques are effective to improve inference speed while maintaining perplexity in text generation He et al. (2021a); Xu et al. (2023). Therefore, inspired by these works for faster NN-LM, we investigate how the approximate nearest neighbor search techniques, such as Product Quantization Jégou et al. (2011) (PQ), Inverted File (IVF) clustering and dimension reduction affect the UE and inference speed of our NN-UE. Description of approximate nearest neighbor search techniques and detailed discussion when each method is individually applied to NN-UE are in Appendix I.
Results of Combination of PQ, IVF and Dimension Reduction
We evaluate the UE performance and inference speed when applying PQ, IVF and dimension reduction are applied. Table 7 shows the results on OntoNotes 5.0 and test sets as ID/OOD, respectively. The detailed discussion when changing the parameters of PQ, clustering and dimension reduction are shown in Appendix I. We can see that ECE and MCE are degraded when PQ, IVF and dimension reduction by PCA are applied simultaneously to NN-UE.121212Distance recomputation does not mitigate this behavior, see Appendix J. On the other hand, our results show that applying them appropriately such as combining PQ with IVF improve inference time with mitigating the degradation in UE performance (The results with the parameters for PQ or IVF can be found in Appendix I.1 or I.2). To deepen our understanding of the above changes in the behavior of the uncertainty performance due to appling of approximate NN search techniques or dimension reduction in NN-UE, we calculated the coverage that how much the indices obtained when using the default exhaustive search are covered when applying PQ, clustering, and dimension reduction sequentially.
Table 8 shows the coverages on OntoNotes 5.0 and as ID/OOD settings, respectively. We can see that applying PQ, clustering, and PCA simultaneously hardly covers any of the indices from the default NN-UE. It is assumed that applying PQ and PCA in the same time leads to coarse distance computation in a single subvector, which would correspondingly degrade the UE performance in NN-UE. Actually, the experimental results in Table 17 in Appendix I.3 suggest that excessive dimension reduction in distance computation could have a negative impact on the UE performance. On the other hand, if combined with PQ and IVF, or applied PCA individually, some of the ground-truth nearest neighbor examples still exist.
8 Conclusion
In this paper, we proposed NN-UE, which estimates uncertainty by using the distance to neighbors and labels of neighbors. The experimental results showed that our method showed higher UE performance than existing UE methods in SA, NLI and NER. Furthermore, our analysis of correctly and incorrectly predicted instances suggests that the improvement in NN-UE is largely due to the reduction in confidence on incorrect instances. In addition, we investigated the effects of efficient neighbor search techniques in NN-UE to address the degradation of the inference speed in token-level tasks such as NER. As a result, we found that product quantization, clustering, or dimension reduction improves inference speed without degrading the UE much more, unless combining all of them simultaneously.
9 Limitations
In this study, we focused only on the classification-based tasks. On the other hand, taking advantage of the recent growth of Large Language Models, UE in text generation is also attracting attention Yoshikawa and Okazaki (2023); Fadeeva et al. (2023); Lin et al. (2024). Therefore, to investigate the effectiveness of NN-UE in text generation tasks is an interesting direction for future research. Not only that, our proposed method is applicable to more tasks such as image classification.
Furthermore, although NN-UE only used the representation of the last layer of the base model, exploring for an appropriate representation for UE is a future challenge. Also, to investigate the relationship between the representation quality and in- and out-of-domain UE performance when using smaller pretrained encoders than DeBERTa, such as BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) is an interesting direction.
Finally, we used ECE and MCE to measure calibration performance. On the other hand, it may be more appropriate to use other metrics to measure calibration performance when the dataset with multiple annotations including human disagreement is available Baan et al. (2022), where it may be similar to the label disagreement in similar output representations. In Section 7.2, we showed that our NN-UE with the label term improves in the direction we expected: it reduces confidence much more when the predictions are inaccurate. However, measuring calibration performance on a variety of data with multiple annotations may provide a more interesting insight into the behavior of our proposed method.
Ethical Considerations
In this study, we used existing datasets that have cleared ethical issues following policies of published conferences. Therefore, they do not introduce any ethical problems. On the other hand, we have an ethical consideration about UE. Specifically, decision support systems with machine learning algorithms do not necessarily have a positive effect on performance. Jacobs et al. (2021) showed that collaboration with machine learning models does not significantly improve clinician’s treatment selection performance, and that performance is significantly degraded due to the presentation of incorrect recommendations. This problem is expected to remain even if UE methods are applied to machine learning models. In addition, introducing UE methods could conversely lead humans to give overconfidence in machine learning models, resulting in performance degradation.
Acknowledgements
The authors also acknowledge the Nara Institute of Science and Technology’s HPC resources made available for conducting the research reported in this paper.
Appendix A Dataset Statistics
The dataset statistics in our study is shown in Table 9.
| Tasks | Datasets | Train | Val | Test | |
| SA | IMDb | 2 | 25,000 | 12,500 | 12,500 |
| Yelp | 2 | - | - | 19,000 | |
| NLI | MNLI | 3 | 392,702 | 4,907 | 4,908 |
| SNLI | 3 | - | - | 9,824 | |
| NER | OntoNotes 5.0 () | 37 | 10,683 | 1,295 | 1,357 |
| OntoNotes 5.0 () | 37 | - | - | 2,327 | |
| OntoNotes 5.0 () | 37 | - | - | 1,366 |
Appendix B Training Settings for Density Estimator in Density Softmax
In Density Softmax Bui and Liu (2024), we use RealNVP Dinh et al. (2017) as the density estimator, which has two coupling structures. Table 10 shows the hyperparameters for training RealNVP as the density estimator in Density Softmax.
| Hyperparameters | Values |
| learning rate | 1e-4 |
| optimizer | AdamW Loshchilov and Hutter (2019) |
| early stopping patient | 5 |
| number of coupling layers | 4 |
| hidden units | 16 |
Appendix C Details of Baselines
Softmax Response (SR)
is a trivial baseline, which treats the maximum score from output of the base model’s softmax layer as the confidence Cordella et al. (1995).
Temperature Scaling (TS)
is a calibration technique by which the logits are divided by a temperature parameter before applying the softmax function Guo et al. (2017). We optimized by L-BFGS on validation set loss.
Label Smoothing (LS)
MC Dropout
is an UE technique by times stochastic inferences with activating dropout Gal and Ghahramani (2016). In our experiments, we set for all evaluations, and the dropout rate is 0.1.
Spectral-Normalized Gaussian Process (SNGP)
uses spectral normalization of the weights for distance-preserving representation and Gaussian Processes in the output layer for estimating uncertainty Liu et al. (2020).
Posterior Networks (PN)
Mahalanobis Distance with Spectral-Normalized Network (MDSN)
is a Mahalanobis distance based UE method that benefits from by spectral normalization of the weights Vazhentsev et al. (2022), similar to SNGP.
E-NER
applies EDL framework for NER by introducing uncertainty-guided loss terms Zhang et al. (2023).
Appendix D Details of Evaluation Metrics
Expected Calibration Error (ECE)
ECE Naeini et al. (2015) quantifies the difference between the accuracy and confidence of a model. Formally, ECE is expressed as:
where is the number of confidence interval bins, denotes the set of examples with predicted confidence scores in the -th bin, is the total number of examples, is the accuracy of the model on the examples in , and is the average confidence of the model on the examples in . In this study, we use .
Maximum Calibration Error (MCE)
MCE, as detailed by Naeini et al. (2015) measures the maximum difference between the model’s accuracy and the confidence across variousb confidence levels. MCE is defined as:
A lower MCE means that there is a small risk that the confidence of the model’s prediction will deviate greatly from the actual correct answer. In this study, we use , same as ECE.
Area Under the Risk-Coverage curve (AURC)
The AURC is the area of the risk-coverage curve when the confidence levels of the forecasts corresponding to the data points are sorted in descending order. The larger the area, the lower the error rate corresponding to a higher confidence level, which means that the output confidence level is more appropriate. Formally, AURC is defined as:
where returns 1 if the prediction is wrong and 0 otherwise.
Excess-Area Under the Risk-Coverage curve (E-AURC)
E-AURC Geifman et al. (2019) is a measure of the AURC score normalized by the smallest risk-coverage curve area , where is the error rate of the model. The reason for normalizing the AURC is that the AURC depends on the predictive performance of the model and allows for performance comparisons of confidence across different models and training methods. E-AURC is defined as:
| E-AURC | (9) |
E-AURC scores are reported with multiplying by 1,000 due to visibility.
Appendix E Additional Results on the Brier score
The Brier score is a widely used metric in UE community for evaluating the probabilistic predictions. The metric measures the mean squared difference between the predicted probability assigned to the predicted label and the actual outcome. This evaluation serves as a holistic assessment of model performance, reflecting both fit and calibration, in the following formula:
| Brier score | (10) |
where is the predicted probability assigned to the prediction, and is the actual outcome. Table 11 shows the results on the Brier score. These results indicate NN-UE improves calibration performance more prominently than other methods while maintaining prediction performance.
| Methods | SA | NLI | ||
| IMDb | Yelp | MNLI | SNLI | |
| SR | 5.000.27 | 5.830.98 | 9.500.40 | 11.020.41 |
| TS | 5.090.42 | 6.671.36 | 8.310.25 | 9.600.21 |
| LS | 4.640.23 | 5.160.92 | 8.730.23 | 10.180.17 |
| MC Dropout | 4.880.21 | 5.450.55 | 9.330.36 | 11.000.28 |
| SNGP | 4.780.15 | 5.990.39 | 12.255.38 | 13.454.57 |
| PN | 10.310.28 | 11.160.22 | 20.760.32 | 21.110.42 |
| Density Softmax | 4.820.18 | 6.050.38 | 9.600.34 | 11.280.41 |
| DAC | 4.440.33 | 5.440.71 | 8.210.25 | 9.550.35 |
| kNN-UE (w/o label) | 4.370.16 | 5.100.12 | 8.150.15 | 9.520.32 |
| kNN-UE | 4.210.14 | 5.020.42 | 8.070.18 | 9.440.28 |
Appendix F The impact of NN-UE with label term in NER on E-AURC
NER tasks are often in label imbalanced settings, where the "O" label is typically much more than other entity-related labels. Additionally, in the OntoNotes 5.0 dataset, the number of labels is 37, as shown in Table 9, which is significantly higher than in SA and NLI tasks. As a result, compared to SA and NLI, neighbor labels will contain much more different labels from the predicted label. The presence of many other labels in the neighbors that are different from the predicted label can lead to excessively low confidence in NN-UE using the label term, even though the prediction is correct because in the label term becomes lower in NER. The impact of that bias for calibration errors, such as ECE and MCE, will be limited. However, low confidence in accurate prediction reduces the coverage much in E-AURC, leading to a degradation in E-AURC.
For example, assume that in a test data set of 6 examples for 3 classes classification, the predictions for the first 3 examples are incorrect and the latter 3 examples are correct. In case A, we assume that prediction confidences are [[0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.02, 0.02, 0.96], [0.01, 0.01, 0.98]]. In case B, we assume that prediction confidences are [[0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.275, 0.275, 0.45], [0.02, 0.02, 0.96], [0.01, 0.01, 0.98]]. In these settings, the ECE and E-AURC for each case are shown in Table 12. These scores indicate that E-AURC is strongly penalized when the confidence in a correct prediction is lower than the confidence in an incorrect prediction.
| Case | ECE (↓) | E-AURC (↓) |
| A | 17.67 | 18.80 |
| B | 16.83 | 121.57 |
Appendix G Inference Time Full Results
We show the inference time full results on out-of-domain test sets in Table 13.
| Methods | SNLI | OntoNotes 5.0 |
| SR | 21.590.76 | 5.750.27 |
| TS | 21.640.07 | 5.790.17 |
| LS | 21.700.07 | 5.800.19 |
| MC Dropout | 396.861.10 | 101.980.83 |
| SNGP | 24.590.08 | - |
| PN | 23.260.05 | - |
| MDSN | 23.390.85 | - |
| E-NER | - | 5.780.61 |
| Density Softmax | 22.020.05 | 6.020.07 |
| DAC | 2346.6236.06 | 326.001.41 |
| NN-UE (w/o label) | 23.020.04 | 10.360.21 |
| NN-UE | 23.070.05 | 10.480.12 |
Appendix H Inference Time Results When Changing -
To estimate whether the inference time changes significantly when changing Top- in kNN search, we investigated the inference time when changing on the IMDb test set. Table 14 shows that the inference time remains almost the same when changing in the range of 8 to 128.
| Methods | Inference time [s] |
| SR | 121.560.12 |
| NN-UE (=8) | 128.980.11 |
| NN-UE (=16) | 128.970.13 |
| NN-UE (=32) | 128.540.16 |
| NN-UE (=64) | 129.160.16 |
| NN-UE (=128) | 128.390.20 |
Appendix I Each Result of Product Quantization, Clustering, and Dimension Reduction
I.1 Product Quantization
(PQ) Jégou et al. (2011) is a data compression technique based on vector quantization. In PQ, a -dimensional representation is divided into subvectors and quantized by performing -means clustering on the vectors in each subspace. Vector quantization can significantly reduce the amount of memory occupied by vectors.131313For example, raw datastore in kNN-UE is 636MB on OntoNotes 5.0 , but PQ reduces it to 10MB. In addition, by calculating the distance between compressed PQ codes, we can efficiently calculate the estimated value of the original Euclidean distance.
We evaluated UE performance and inference time when the number of clusters in the codebook was fixed at 32, and the number of subvectors was changed to (In Table 7 and 8, PQ was performed with ).
Table 15 shows the UE performance and inference time results in different . In ECE and E-AURC, there are almost no degradation in UE performance due to PQ. On the other hand, in MCE in ID setting, the UE performance consistently degrades. Furthermore, compared to NN-UE among different , the larger , the better the UE performance tends to improve, but the inference time increases.
The larger is, the more time is required for inference but the UE performance improves. We assumed that these results are derived from the decrease in quantization error over the vector with PQ with larger because each subvector is divided into smaller subspaces and the quantization is performed for each subspace. On the other hand, an increase in requires additional distance computations etc., then more inference time.
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| OntoNotes 5.0 (In-domain) | ||||
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 2.490.08 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 4.940.10 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 |
| NN-UE () | 1.900.27 | 31.1811.17 | 20.161.12 | 3.270.06 |
| NN-UE () | 1.960.31 | 31.3318.74 | 20.231.27 | 3.320.05 |
| NN-UE () | 1.880.34 | 31.0616.36 | 20.161.23 | 4.110.11 |
| OntoNotes 5.0 (Out-of-domain) | ||||
| SR | 17.050.69 | 37.063.13 | 81.494.17 | 5.750.27 |
| NN-UE (w/o label) | 8.780.62 | 24.911.81 | 70.104.03 | 10.360.21 |
| NN-UE | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| NN-UE () | 7.660.48 | 17.073.81 | 74.475.53 | 7.220.19 |
| NN-UE () | 7.570.45 | 16.432.73 | 74.385.36 | 7.230.16 |
| NN-UE () | 7.570.44 | 16.382.66 | 74.355.49 | 8.900.18 |
I.2 Clustering
The original NN-LM uses an inverted file index (IVF) technique that speeds up the search by dividing the representation into clusters by -means and searching for neighbors based on centroids. In this study, we evaluate the UE performance and inference speed when the number of clusters . In this study, we evaluate the UE performance and inference speed when the number of clusters and applying PQ with are fixed and the number of cluster centroids to search changes (In Table 7 and 8, IVF was performed with ).
Table 16 shows the performance of UE when changing in ID and OOD settings using OntoNotes 5.0. In ECE, scores are slightly reduced for ID, but only slightly worse for OOD; MCE also shows degradation for ID but little for OOD, and even improves when ; E-AURC shows almost no change in scores when is changed for both ID and OOD. In terms of inference time, the larger , the longer it takes. We derive the improvement in MCE when increasing in ID setting from the fact that more clusters are targeted, making it possible to cover ground-truth nearest neighbor examples. On the other hand, the tendency of slight decrease when increasing in OOD setting may comes from the reliability of the vector, similar to the discussion in Section 6.3.
In addition, Taken together with the results in Table 7 in Section 7.3, we can see that the degradation of the UE performance can be mitigated with improvement latency when applying PQ and IVF with lower , compared to applying PQ, IVF and PCA simultaneously.
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| OntoNotes 5.0 (In-domain) | ||||
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 2.490.08 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 4.940.10 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 |
| NN-UE () | 1.820.28 | 30.1816.77 | 20.141.21 | 2.840.08 |
| NN-UE () | 1.860.25 | 29.4816.91 | 20.131.21 | 3.110.03 |
| NN-UE () | 1.920.31 | 28.5511.24 | 20.131.22 | 3.310.06 |
| NN-UE () | 1.830.28 | 27.009.43 | 20.141.21 | 3.710.06 |
| OntoNotes 5.0 (Out-of-domain) | ||||
| SR | 17.050.69 | 37.063.13 | 81.494.17 | 5.750.27 |
| NN-UE (w/o label) | 8.780.62 | 24.911.81 | 70.104.03 | 10.360.21 |
| NN-UE | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| NN-UE () | 7.520.41 | 16.011.92 | 74.335.37 | 6.090.28 |
| NN-UE () | 7.560.36 | 16.933.38 | 74.315.39 | 6.650.17 |
| NN-UE () | 7.600.41 | 17.122.35 | 74.345.35 | 7.330.21 |
| NN-UE () | 7.530.40 | 17.282.45 | 74.335.37 | 7.890.12 |
I.3 Dimension Reduction
In general, Transformer-based models such as PLM have high-dimensional token representations. In high-dimensional spaces, nearest neighbor search often suffer from the curse of dimensionality. To reduce this problem, we apply dimension reduction to NN-UE similar to He et al. (2021a). In this study, we use Principal Component Analysis (PCA) as a dimension reduction algorithm to reduce the dimension of the datastore representations and the query representation (In Table 7 and 8, PCA was performed with ). As shown in Table 17, the UE performance depends on the number of target dimensions, and the performance degrades when or . On the other hand, the performance in is almost the same as default NN-UE. This suggests that excessive dimension reduction in distance computation to extract nearest examples by NN search could have a negative impact on the UE performance.
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| OntoNotes 5.0 (In-domain) | ||||
| SR | 7.790.53 | 50.0724.15 | 21.901.31 | 2.490.08 |
| NN-UE (w/o label) | 3.370.71 | 33.153.65 | 17.630.66 | 4.940.10 |
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 |
| NN-UE () | 1.890.37 | 31.0114.35 | 20.061.25 | 3.240.08 |
| NN-UE () | 1.800.36 | 27.8513.80 | 20.131.29 | 3.410.10 |
| NN-UE () | 1.800.40 | 26.2312.61 | 20.131.28 | 3.850.06 |
| OntoNotes 5.0 (Out-of-domain) | ||||
| SR | 17.050.69 | 37.063.13 | 81.494.17 | 5.750.27 |
| NN-UE (w/o label) | 8.780.62 | 24.911.81 | 70.104.03 | 10.360.21 |
| NN-UE | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| NN-UE () | 7.480.41 | 16.202.75 | 74.335.49 | 7.370.26 |
| NN-UE () | 7.540.45 | 16.422.73 | 74.305.44 | 7.750.24 |
| NN-UE () | 7.560.43 | 16.132.59 | 74.265.40 | 8.510.46 |
Appendix J Distance Recomputation for NN-UE
When using efficient NN search techniques in Section 7.3, we use approximate distances to compute Eq. 4. Although we can get raw vectors by using the example indices obtained from approximate nearest neighbor search and compute accurate distance, in NN-LM this has been shown to lead to performance gains and latency degradation He et al. (2021a). We measure the UE performance and inference speed when PQ, clustering, and dimension reduction are applied simultaneously and re-computing accurate distances, reported in Table 18. These results show that the UE performance does not improve except for MCE in the ID setting, and the latency is about 5-7x slower when reading raw vectors from the datastore and re-computing distances. Moreover, these results suggest that exact distance computation for examples that are not actually nearest neighbors are not very effective in NN-UE.
| Methods | ECE (↓) | MCE (↓) | E-AURC (↓) | time [s] |
| OntoNotes 5.0 (In-domain) | ||||
| NN-UE | 1.780.32 | 26.0213.72 | 20.141.27 | 4.990.07 |
| NN-UE (Approx.) | 2.140.37 | 33.5210.84 | 20.121.26 | 2.870.04 |
| NN-UE (Recomp.) | 2.350.44 | 30.477.50 | 20.161.17 | 16.240.77 |
| OntoNotes 5.0 (Out-of-domain) | ||||
| NN-UE | 7.500.42 | 16.532.61 | 74.275.43 | 10.480.12 |
| NN-UE (Approx.) | 8.080.53 | 24.035.46 | 74.505.42 | 6.200.20 |
| NN-UE (Recomp.) | 8.300.51 | 25.675.26 | 74.585.53 | 34.220.78 |
Appendix K Licenses of Datasets, Tools and Models
Datasets
The IMDb movie dataset can be used for research purposes as described in https://developer.imdb.com/non-commercial-datasets/. Yelp dataset can be used for academic purposes as described in https://s3-media0.fl.yelpcdn.com/assets/srv0/engineering_pages/f64cb2d3efcc/assets/vendor/Dataset_User_Agreement.pdf. The MNLI dataset is licensed for research purposes as described in Williams et al. (2018). The SNLI dataset can be used for research purposes as described in https://nlp.stanford.edu/projects/snli/. OntoNotes 5.0 dataset can be used for research purposes as described in https://catalog.ldc.upenn.edu/LDC2013T19.
Tools
is licensed by Apache-2.0. is MIT-licensed.
Models
and from Huggingface model checkpoints are MIT-licensed.